2x+3y=1
-3x+y=15



5r-s=5
-4r+5s=17



x-3y=0
3x+y=7



0.5x-2y=17
2x+y=104
please help solve these linear equations by using sub.

Answer by RAY100(1637)

problem 1
1) 2x+3y=1
2)-3x+y=15,,,,,y=(15+3x)
subst into (1)
2x+3(15+3x)=1
2x+45+9x=1
11x=-44
x=-4
y=15+3x=15+3(-4)=15-12=3
check by subst in (1) &(2) ok
problem 2
1) 5r-s=5,,,,,s=-5+5r=5r-5
2) -4r+5s=17
subst into (2)
-4r+5(5r-5)=17
-4r+25r-25=17
21r=42
r=2
S=5r-5=5(2)-5=5
check in (1) &(2) ok
problem 3
1)x-3y=0,,,,x=3y
2)3x+y=7
subst into (2)
3(3y) +y=7
9y+y=7
10y=7
y=7/10
x=3y=3(7/10)=21/10
check in (1)&(2) ok
problem 4
1) 1/2 (x) -2y=17
2) 2x+y=104,,,,,y=104-2x
subst into (1)
1/2(x) -2(104-2x)=17
1/2(x) -208+4x=17
4.5x=225
x=50
y=104-2x
y=104-2(50)=104-100=4
check in (1) &(2) ok